OpenStudy (anonymous):

Write an equation so y is a function of x: 8x - 2y = 16

jimthompson5910 (jim_thompson5910):

The goal is to solve for y. To do so, you subtract 8x from both sides. Then you divide both sides by -2 to fully isolate y. \[\Large 8x - 2y = 16\] \[\Large -2y = 16-8x\] \[\Large -2y = -8x+16\] \[\Large y = \frac{-8x+16}{-2}\] \[\Large y = \frac{-8x}{-2}+\frac{16}{-2}\] \[\Large y = 4x-8\]